OpenStudy (anonymous):

Help please... Estimate the number of atoms in 1 cm3 of a solid. (Note that the diameter of an atom is about 10 ^-10 m

OpenStudy (xishem):

It really depends on the way that the atoms are packed. Does the problem give any other indication to the structure of the unit cell?

OpenStudy (anonymous):

no :(

OpenStudy (xishem):

Ok. We will assume that is is packed in a cubic manner, which is basically, like this in all 3 dimensions:

OpenStudy (shane_b):

It certainly depends on how they are packed but since it's just an *estimate* I'd go with something like this: \[\large V_{sphere}=\frac{4}{3}\pi r^3=\frac{4}{3}\pi*(\frac{10^{-10}}{2}m)^3=5.24 x 10 ^{-31}m^3\]\[\large 1cm^3=1x10^-6m^3\]Dividing the volume of a single atom into the available volume you'd get: \[\large \frac{1x10^{-6}m^3}{5.24 x 10 ^{-31}m^3}=1.91x10^{24}atoms\]

OpenStudy (xishem):

I took a different route to estimation, and it still ended up being within a factor of 2, so that just shows you how much the answer may vary based on what packing method you choose.

OpenStudy (xishem):

My method is wrong, however, because I assumed that the atoms were cubic not spherical :P. Shane_B's approximation is closer.

OpenStudy (shane_b):

With that many atoms....what's a factor of 2 really :) If the teacher doesn't like your estimate, tell her to count them :P

OpenStudy (xishem):

Exactly :P.

OpenStudy (anonymous):

Thank you! I really appreciate the help. Wish my teacher was like you two...

OpenStudy (shane_b):

you're welcome :)